A collection of 21 nickels, dimes, and quarters is worth $2.50. There are five fewer dimes than nickels. How many quarters are there in the collection?
Answer by jorel1380(3719)

x+y+z=21
x-5=y
x+(x-5)+z=21
2x+z=26
5x+10(x-5)+25z=250
5x+10x-50+25z=250
15x+25z=300
50x+25z=650
35x=350
x=10
y=5
z=6
There are 10 nickels, 5 dimes, and 6 quarters in the collection..